OpenStudy (anonymous):

whats the mean, median, and mode of the numbers 85,83,85,82

jimthompson5910 (jim_thompson5910):

mean = add up the numbers, then divide that result by 4

jimthompson5910 (jim_thompson5910):

median = middle most number so you sort the numbers in order and pick out the middle most number (in this case, you have to average the two middle most numbers to get the median)

jimthompson5910 (jim_thompson5910):

mode = most frequent number so you just write down the number that shows up the most

OpenStudy (anonymous):

how bout the range ?

jimthompson5910 (jim_thompson5910):

range = (largest number) - (smallest number)

OpenStudy (anonymous):

how do you do median again ? what do you do to get it ?

jimthompson5910 (jim_thompson5910):

ex: say you had the numbers 3, 1, 2, 5, 4 you would sort them to get this 1, 2, 3, 4, 5 the middle most number is 3 so the median is 3

OpenStudy (anonymous):

but what if you have 2 middle numbers?

jimthompson5910 (jim_thompson5910):

now say that you only have 1, 2, 3, 4 the middle most numbers are 2 and 3 average them to get (2+3)/2 = 5/2 = 2.5

jimthompson5910 (jim_thompson5910):

so the median of 1,2,3,4 is 2.5

OpenStudy (anonymous):

thank you !!! so much !

jimthompson5910 (jim_thompson5910):

you're welcome